A one-year discount bond issued by X has a payout of $1,250 and today's price is $1,189. A one-year discount bond issued by Y has a payout of $382 and today's price is $354. Then the bond issued by X has a ____ yield than the bond issued by Y, and this could be because X has a ____ default risk.

Group of answer choices

a) higher; lower

b) lower; lower

c) lower; higher

d) higher; higher
Answer #1

Answer is b) lower, lower

Explanation:

Yield is calculated by formula: payouts/price-1 (since it's a 1-year zero coupon bond)

yield for X =(1250/1189)-1 = .05 = 5%

yeild for Y= (382/354)-1 = .079 = 8%

Since higher yields indicate that risk associated with bond is higher(because you need to increase the yields of a bond to make it more attractive for people to buy), so answer is (b)